For a particular reaction at 124.8 °C, Δ?=−841.80 kJ/mol , and Δ?=646.52 J/(mol⋅K) . Calculate ΔG for this reaction at 11.9 °C.
Answer #1

Hera given 8418KJmol at Tempenalunu b124-8c A G= - 1248(124.8+ 273) K= 397.8 K A S = 646.5L J mol k A G = AH-TAS AH= 4 G+ T 4